Kumar Co. is attempting to predict its maintenance costs more accurately. Maintenance costs are a mixed cost. Maintenance costs and machine hours for the first four months of the year are as follows:

Month Maintenance Costs Machine Hours
Jan. $8,430 1,320
Feb. 7,620 1,190
Mar. 8,810 1,430
Apr. 7,580 1,130


Required:

Using the high-low method, calculate unit variable cost and monthly fixed costs.
Answer #1

Calculation of Variable Cost per Unit:

Variable Cost per Unit = (y2 - y1) / (x2 - x1)

where y2 = the total cost at highest level of activity = 8,810
y1 = the total cost at lowest level of activity = 7,580
x2 = the number of labor hours at highest level of activity = 1,430
x1 = the number of labor hours at lowest level of activity = 1,130

So, Variable cost per unit = (8,810 - 7,580) / (1,430 - 1,130)

= 1,230 / 300 = $4.10 Ans.

Calculation of Monthly Fixed Cost:

Monthly Fixed Cost = y2 − bx2 = y1 − bx1

where y2 = the total cost at highest level of activity = 8,810
   y1 = the total cost at lowest level of activity = 7,580
   x2 = the number of labor hours at highest level of activity = 1,430
  x1 = the number of labor hours at lowest level of activity = 1,130

b = Variable cost per unit

So, Monthly fixed cost = 8,810 - (4.10 * 1,430) = 7,580 - (4.10 * 1,130) =$2,947 Ans.